OpenStudy (anonymous):

In a sample of seawater taken from an oil spill region, it is found that an oil layer 4 cm thick floats on 55cm of water. If the density of the oil is 0.75 x 10^3 kg/m3, what is the absolute pressure on the bottom of the container?

OpenStudy (anonymous):

sum two pressures based on p = rho g h p1 = (oil density) (0.04m) (9.8 N/kg) p2 = (water density) (0.55 m) (9.8 N/kg) p = p1 + p2 in N/m^2 also called Pa

OpenStudy (anonymous):

thank you again...what's the formula for gauge pressure? absolute pressure? :))

OpenStudy (anonymous):

absolute pressure = atmospheric pressure plus gauge pressure psia = atm + psig

OpenStudy (anonymous):

thank you ! ^_^